OpenStudy (anonymous):

WHAT IS 7 AND 1/3 PLUS 8 AND 3/4 ESTIMATED????

OpenStudy (anonymous):

7 + 1/3 = 21/3 + 1/3 = 22/3 8 + 3/4 = 32/4 + 3/4 = 35/4 So, 22/3 plus 35/4 = 4(22)/12 + 3(35)/12 = (88+105)/12 = 193/12 = 16 and 1/12